In a random sample of 144 voters it was found that 90 were in favor of a bond issue. Test the
hypothesis that opinion is equally divided on this issue. Let α = 0.01. Hint, voters are either pro or con
regarding the bond issue and the random sample is small relative to the population of voters, so voters
are independent of each other regarding attitude toward the bond issue.

Answer by rothauserc(4718)

Ho: X = 0.50
:
H1: X > 0.50
:
sample mean proportion = 90/144 = 0.625
:
standard error(SE) = square root( 0.625 * (1-0.625) / 144) = 0.0403
:
Note we use the sample proportion because we do not know the population proportion
:
test statistic = (0.625 - 0.500)/0.0403 = 3.10
:
alpha(a) = 0.01
:
critical probability(p*) = 1 - (a/2) = 0.995
:
the critical value(z-score) associated with p* is 2.57
:
Since the test statistic(3.10) does fall within the critical region(2.57), we reject the null hypothesis.
: